### v2

# Who is better at math, Jenny or Jingzhen? Uncovering Stereotypes in Large Language Models

Abstract
Large language models (LLMs) have been shown to propagate and amplify harmful stereotypes, particularly those that disproportionately affect marginalised communities. To understand the effect of these stereotypes more comprehensively, we introduce GlobalBias, a dataset of 876k sentences incorporating 40 distinct gender-by-ethnicity groups alongside descriptors typically used in bias literature, which enables us to study a broad set of stereotypes from around the world. We use GlobalBias to directly probe a suite of LMs via perplexity, which we use as a proxy to determine how certain stereotypes are represented in the model’s internal representations. Following this, we generate character profiles based on given names and evaluate the prevalence of stereotypes in model outputs. We find that the demographic groups associated with various stereotypes remain consistent across model likelihoods and model outputs. Furthermore, larger models consistently display higher levels of stereotypical outputs, even when explicitly instructed not to.
1 Introduction
LLMs are increasingly used for tasks that span areas of concern for bias and fairness Anthis et al. (2024), such as user discrimination in recommendations Xu et al. (2023). Despite the obvious need for ethical frameworks around these models, these are mostly lacking or incomplete, and make research into fairness and bias essential for supporting public confidence in the use of generative AI. While bias is often defined in ambiguous and conflicting ways Blodgett et al. (2020), in this paper we focus on representational harms, defined by Crawford (2017) as harms that “occur when systems reinforce the subordination of some groups along the lines of identity,” specifically harms caused by stereotyping.
Existing research on stereotypes in LMs is limited, and predominantly focuses on African American and White groups Jiang and Fellbaum (2020); May et al. (2019), or a subset of US census groups, often with Middle Eastern added Guo and Caliskan (2020); Cao et al. (2022); Kirk et al. (2021); Cheng et al. (2023). Furthermore, datasets that seek to expand the coverage of bias measures to multiple axes are limited to a fixed set of stereotypes for specific demographic groups Nangia et al. (2020); Nadeem et al. (2021); Parrish et al. (2022). To address these limitations, we focus on incorporating a wide range of ethnicities, allowing us to gain a more international view of the effects of bias. We also highlight the importance of analysis that uses a intersectional lens, where biases compound across a combination of different axes, e.g., gender and ethnicity, to cause unique harms.
Attribute Details
Name Kazuyo
Age 45
Personality Traits Calm, Wise, Observant
Negative Traits Perfectionist, Indecisive, Shy
Hobbies Bonsai gardening, Origami, Tea ceremonies
Occupation Librarian
Height 5.2 ft
Hair Colour Black
Eye Colour Brown
Skin Colour Light
Build Petite
Socioeconomic Status Middle class
Sexual Orientation Asexual
Religion Shinto
Table 1: Example of a character profile generated by Claude 3 Opus for given name Kazuyo (Japanese, F). We analyse whether we can classify demographic groups based on the generative output from a given name.
In Sections 4 and 5, we utilise templates involving stereotypes for 40 groups, defined by both an ethnicity and a gender, e.g. English Female or Chinese Female, along with a descriptor (e.g. ’good at math’) to explore which descriptors are more likely to appear in a sentence with certain given names across different LLMs.
Considering the limitations of using a fixed set of stereotypes and the fact that likelihoods do not always correspond to model outputs Parrish et al. (2022), in Section 6, we take a lexicon-free approach that utilizes the given names in our dataset in a generation task. An example output can be seen in Table 1. We present both quantitative and qualitative analyses of representational harms caused by stereotypical outputs. The results highlight the magnitude of stereotypical bias across both open and closed-sourced LLMs. From this, this work presents the following contributions:
  1. 1.
    the GlobalBias [^1] The dataset and code used to evaluate the models can be found at https://github.com/groovychoons/GlobalBias. dataset for studying harmful stereotypes, which consists of 876,000 sentences for 40 distinct gender-by-ethnicity groups
  2. 2.
    an analysis of which stereotypes are surfaced for each group by a number of LMs, and the extent and nature of harm caused by the these stereotypes, particularly for intersectional groups
  3. 3.
    the finding that larger models have more stereotypical outputs, even when explicitly instructed to avoid stereotypes and clichés
  4. 4.
    the finding that bias stays consistent across model’s internal representation and outputs, contrary to claims in previous work in the field.
2 Background and Related Work
2.1 Impact of Stereotyping
Stereotyping can influence how we perceive ourselves and others, as well as how we behave towards others. For example, Bertrand and Mullainathan (2004) found résumés with White names received 50% more invitations to interview than resumes with Black names. More broadly, Biernat (2003) found that when one judges individual members of stereotyped groups on stereotyped dimensions, one does so with reference to within-category standards, e.g. evaluations of men and women on leadership competence may not be directly comparable, as their meaning is tied to different referents: ‘good’ for a woman does not mean the same thing as ‘good’ for a man. LLMs trained on data that includes stereotypes or LLMs using non-comprehensive systems to mitigate biases can perpetuate discrimination and social inequality in ways that are difficult to detect and address.
2.2 Axes of Analysis
Early work on bias in word embeddings focused on a single dimension, predominantly binary gender Bolukbasi et al. (2016); Zhao et al. (2018b); Ethayarajh et al. (2019), and less frequently, race Caliskan et al. (2017); Garg et al. (2018). Work looking at a single demographic axis often fails to mirror the reality of race and gender being intertwined.
Crenshaw (1989) defines how using a single-axis framework erases Black women’s experience in legal and political contexts, as race discrimination tends to be viewed in terms of gender-privileged Blacks, and gender discrimination focuses on race-privileged women. Crenshaw provides a framework for understanding how different aspects of a person’s social and political identities combine to create different modes of discrimination and privilege, known as intersectionality.
There is a growing body of research in the field of intersectional bias, which starts to investigate the nuance of how race and gender interact Jiang and Fellbaum (2020). There are several measures defined for evaluating intersectional biases, such as the angry black woman stereotype in contextual word embeddings May et al. (2019); Tan and Celis (2019), the contextual word embedding test (CEAT) which also looks at a limited and fixed labelled set of stereotypes Guo and Caliskan (2020), and others Lepori (2020); Cao et al. (2022); Cheng et al. (2023).
2.3 Stereotype Datasets
This paper builds on previous work using stereotypes as a means of exploring characteristics that are observed to be associated with specific demographic groups, reinforcing established social hierarchies Greenwald et al. (1998); Blodgett et al. (2021). A number of datasets have been developed to look specifically at stereotypes, largely focusing on using sentence pairs to compare two demographic groups May et al. (2019), a stereotype and anti-stereotype Zhao et al. (2018a); Nangia et al. (2020); Nadeem et al. (2021), or a question and answer set to compare two groups Parrish et al. (2022). While valuable, these datasets are all situated within a U.S. context Blodgett et al. (2021), and are unsuitable for our context of analysing multiple subgroups across a single stereotype. In contrast, our one-vs-all approach offers more robust statistical power, and reduces the impact of outliers and natural variability in two-group comparisons. In terms of metrics, we posit that perplexity, which Smith et al. (2022) and Smith and Williams (2021) used to compare multiple subgroups in one test, is a suitable method, and thus we develop it further.
2.4 Use of Proper Names
There exists measurable statistical tendencies for names to refer to both gender and race demographics Tzioumis (2018). May et al. (2019) observes that "tests based on given names more often find a significant association than those based on group terms". Therefore, we use given names as a proxy for ethnicity and gender, based on evidence that given names are often used to draw stereotypical conclusions about people by both humans Bertrand and Mullainathan (2003); Dechief and Oreopoulos (2012) and in LLM outputs De-Arteaga et al. (2019); Romanov et al. (2019); Maudslay et al. (2019). Using a range of names for each group intends to mitigate the impact of any single name on the group’s overall results.
3 The Dataset
We propose a new dataset named GlobalBias for studying harmful stereotypes, which consists of sets of 10 proper names spanning 40 groups. A summary of key data can be found in Table 2.
3.1 Proper Names
Our primary objective is to compile a list of diverse demographic groups, alongside representative names for each group. In this section, we discuss how we build such a dataset, of specifically 40 distinct groups, starting from existing labeled resources.
Our seed dataset of names is the Genni + Ethnea dataset Torvik (2018). It contains over 2 million names, each annotated with ethnicity and gender. We first filter $\sim$176,000 unique first names, to include only those with 2 to 14 characters and a male or female gender classification, narrowing our dataset to approximately 35,000 names. We exclusively included names labeled with a binary gender by the Genni model used to label the seed dataset, thereby excluding gender-neutral names. We posit that gender-neutral names do not necessarily represent gender diverse groups in LLMs, and are more often a mixture of male and female stereotypes, though we acknowledge that focusing on binary gender classification fails to represent the diverse spectrum of human self-identification as discussed in Butler (1989) and Kuper et al. (2012).
By utilizing embeddings and clustering techniques, we identify names that an LLM perceives as highly correlated within these groups. We use OpenAI’s (text-embedding-3-large) embeddings for each name and apply Mini Batch K-Means clustering to group the names into clusters. We select ten names per group to prevent one name from having a large impact on results. These ten names are randomly selected from clusters with a high gender and ethnicity agreement, i.e. $>50\%$ names in the same group in a cluster, meaning an LLM is likely to classify the chosen cluster as belonging to that ethnic and gender group. Where ethnicities have an exclusive gender, we select 10 names of the opposite gender with high probability of belonging to that ethnicity for gender balance across the entire dataset.
We select 400 unique first names, namely, the part of a personal name considered to distinguish an individual within a group. It is important to note that while naming conventions vary across the world, these names were gathered in a Western academic context where the first name typically corresponds to the given name.
3.2 Descriptors
Having compiled a suitable list of demographic groups and representative names via clustering, our next step is to obtain a set of suitable descriptors. We will combine these descriptors with the names to construct templates, which will serve the input for a probing exercise to various LMs in our experiments. Let us now discuss how we obtain these descriptors, and the resulting templates we derive from them.
We initially draw on three existing datasets: the HOLISTICBIAS dataset Smith et al. (2022), Ghavami and Peplau (2013) and StereoSet Nadeem et al. (2021). The first, HOLISTICBIAS, is split into 13 demographic axes; we use 11 of these axes (Race/Ethnicity and Nationality are excluded, as the purpose of the experiment is to infer these from the given name). Ghavami and Peplau (2013) provides a labelled dataset of stereotypes from a free-response survey. We extract stereotype terms from StereoSet, which was handcrafted to test a fixed set of stereotypes in LLMs. As a result, our descriptor terms represent a diverse range of potential stereotypes.
Parameter Count
Names 400
Descriptors 730
Templates 3
Sentences 876,000
Demographic Groups 40
Table 2: Summary of GlobalBias dataset statistics.
3.3 Templates
Following previous work from Smith et al. (2022), we construct three templates combining the names compiled in Section 3.1 and the descriptors obtained in Section 3.2. These templates allow us to measure token likelihoods of the descriptors in relation to the given names. These templates combine given names and descriptor terms. Examples of the three templates can be found in Figure 1.
At the end of this process, the GlobalBias dataset is ready: it comprises 876,000 sentences covering 40 distinct gender-by-ethnicity groups created through the combination of proper nouns and descriptors. In the next section, we discuss how we use GlobalBias for evaluating stereotypical behaviour in LMs, and discuss the results.
4 Adjusted Perplexity across Descriptors (APX)
4.1 Perplexity
Perplexity has become an increasingly common evaluation measure when looking at stereotypes in LLMs Smith and Williams (2021); Smith et al. (2022). We use perplexity to determine how stereotypical an LM perceives a sentence to be. The lower the perplexity, the more likely an LM is to generate a sequence of words. For decoder-only LMs such as GPT-2 Radford et al. (2019), we compute the perplexity of a tokenized sentence $\boldsymbol{x}=[x_{1}...x_{m}]$ as:
$\text{PPL}({\boldsymbol{x}})=\exp\left(-\frac{1}{m}\sum_{i=1}^{m}\log{P_{\text{lm}}(x_{i}|{\boldsymbol{x}}_{i-1})}\right)$ (1)
where $P_{\text{lm}}(x|{\boldsymbol{x}})$ is the likelihood of the next token given the preceding tokens.
For masked language models (MLM) such as RoBERTa Liu et al. (2019), pseudo-perplexity Salazar et al. (2020) is used instead, which replaces the likelihood $P$ in Equation 1 by $P_{\text{mask}}(x_{i}|{\boldsymbol{x}}_{\neg i})$, the pseudo-likelihood to predict the masked token $x_{i}$ Wang and Cho (2019). For encoder-decoder LMs such as Flan-UL2 Tay et al. (2023), we compute $P_{\text{lm}}$ on the decoder, which is conditioned by the encoder.
4.2 Defining APX
The use of perplexity in this context can be problematic, due to noise from high-frequency given names during training Kaneko and Bollegala (2021), meaning some ethnic and gender groups will tend toward having higher or lower perplexity scores for all descriptors, regardless of any underlying biases. We account for this by proposing a novel bias evaluation metric, which we name Adjusted Perplexity across Descriptors (APX).
Consider the mean perplexity for an intersectional group of given names $G_{i}$ and a descriptor $D_{j}$, we define their perplexity as $\text{PPL}(G_{i}D_{j})$. We define the Adjusted Perplexity across Descriptors to be:
Mean Group $\text{Mean Group Perplexity}=\sum_{j=1}^{D}\frac{\text{PPL}(G_{i}D_{j})}{|D|}$ (2)
[…]
$\text{APX}(G_{i}D_{j})=\text{PPL}(G_{i}D_{j})\times\frac{\text{Mean Group Perp.}}{\text{Mean Total Perp.}}$ (4)
4.3 Models
In our experiments, we evaluate a suite of seven language models to examine the generalizability of our bias measures across various model sizes and architectures, these are: BERT (google-bert/bert-large-cased; Devlin et al. 2018), RoBERTa (roberta-large; Liu et al. 2019), Flan-UL2 (google/flan-ul2, Tay et al. 2023), GPT-2 (gpt2-xl, Radford et al. 2019), GPT Neo X (EleutherAI/gpt-neox-20b; Black et al. 2022), OPT (facebook/opt-30b; Zhang et al. 2022) and Llama 3 (meta-llama/Meta-Llama-3-8B; AI@Meta 2024).
4.4 Validating APX
We measure perplexity and APX on a subset of GlobalBias of 36,960 sentences, composed of 3 templates, 280 unique names, and 44 labeled descriptors, and compare APX to the perplexity metric for classification accuracy and mean reciprocal rank on a range of models. Human participants provide this validation set of racial stereotypes with ground truth information in prior work Ghavami and Peplau (2013). The experiment uses 11 stereotypes for 4 groups, removing any duplicates that appear across multiple groups, for example, ’intelligent’ is associated with both Asian American and White groups.
Two inherent limitations were identified in the dataset. Due to the dataset’s categorization framework of five distinct racial categories, we combined our diverse ethnicities within these predefined categories, eliminating 6 out of 20 ethnicities. The primary objective of this experiment was to validate the APX measure, the full set of ethnicities is explored in more detail in the next experiment. Furthermore, it’s worth noting that the specific focus of African American stereotypes did not correspond directly with given names for any of the ethnic groups under examination, rendering it unsuitable for inclusion within this context.
We take the average of the 10 names per group for each template, and then take the normalised average of the three templates in order to obtain a robust bias score for each gender-by-ethnicity group for each descriptor. To calculate one-vs-all classification accuracy, we take the group with the minimum bias score to be the most biased group. The accuracy shows how often the group with the minimum bias score for each descriptor matches the target group. This methodology enables comparison across masked, encoder-decoder, and decoder-only language models. Despite the variations in how perplexity is calculated for each model type, using the lowest perplexity value from four ethnicity groups ensures the results are generalizable across different model architectures.
Model Acc. (PPL) Acc. (APX)
BERT 38.6% 38.6%
RoBERTa 45.5% 50.0%
Flan-UL2 36.4% 36.4%
GPT-2 31.8% 50.0%
GPT-NeoX 25.0% 38.6%
OPT 36.4% 43.2%
Llama 3 31.8% 50.0%
Table 3: Classification accuracy in a 4 class stereotype classification task. We show the accuracy when using the perplexity and APX metrics for 7 models. Classification accuracy represents how often the group with the minimum bias score for each descriptor matches the target group.
Model MRR (PPL) MRR (APX)
BERT 58.1% 63.6%
RoBERTa 56.6% 69.1%
Flan-UL2 59.3% 62.9%
GPT-2 54.2% 66.5%
GPT-NeoX 54.5% 59.1%
OPT 55.7% 66.1%
Llama 3 58.9% 70.3%
Table 4: Mean Reciprocal Rank in a 4 class stereotype classification task.
Table 3 shows the classification accuracy when using perplexity and APX for the labelled stereotype dataset. We can see that in 5 out of 7 models, the use of APX improves performance, by an average of 12.26%. In addition, we measure Mean Reciprocal Rank (MRR) for each of the 44 descriptors, by ranking the perplexities and APX of the 4 ethnic groups. This allows us to investigate cases where a group may have the second lowest perplexity, which works well for descriptors that may be stereotypes for multiple groups, such as ’family-oriented’ or ’religious’. The results in Table 4 show that using APX improves MRR across all models, with an average improvement of 8.61%.
Our experimental results show that the proposed evaluation measure, APX, outperforms perplexity in classification tasks when assessed using both accuracy and MRR. Thus, APX proves to be a more effective metric for measuring biases in language models. We use APX in the next section to investigate a wider set of demographic groups and stereotypes.
5 Stereotypes via APX
We propose a statistically robust methodology to identify the demographic groups associated with the 730 descriptors in GlobalBias. We calculate the APX for the 876,000 sentences in the dataset. As described in the previous section, we compute the average of the 10 names per group for each template, and take the normalised average of the three templates to obtain a bias score for each gender-by-ethnicity group for each descriptor. Once we have the bias scores for each of the 40 groups, we identify any groups with a 1% one-tailed significance level, as shown in Figure 1. Our methodology can be applied to any descriptor and extended to additional gender-by-ethnicity groups and demographic axes in future.
Figure 1: An overview of our methodology using the example descriptor good at math. We compute the normalised average of APX for 10 names for each template, followed by the average over 3 templates to calculate a bias score. Gender-by-ethnicity groups with a 1% statistical significance (noted by the orange line) are considered to be associated with that descriptor, i.e. Chinese Female with good at math.
5.1 Overview
To ensure consistency and enable comparison across the three experiments detailed in Sections 4, 5, and 6, we use Llama 3 as a case study. We present a full table of results in Appendix A, and a smaller, selected set of descriptors in Table 5, which we refer to in this section. These tables show the descriptors associated with each gender-by-ethnicity group in the Llama 3 8B model.
Overall, we observe the resurfacing of multiple stereotypes noted in other studies, such as associating Arabs with being Muslim and terrorists Chang and Kleiner (2003); Corbin (2017), characterizing Japanese women as shy and cute Zheng (2016); Azhar et al. (2021), and depicting Hispanic males as macho Ghavami and Peplau (2013). Among the 730 descriptors analyzed, 147 (20.1%) demonstrated statistically significant results. This indicates that a substantial portion of descriptors in GlobalBias did not exhibit significant bias towards any specific demographic group. In the following subsections, we discuss the harmful implications of some of the stereotypes uncovered.
Group Selected Descriptors
Arab, F Muslim, refugee
Arab, M extremist, Muslim, terrorist
Chinese, F good at math, quiet, very smart
Hispanic, M macho
Japanese, F always cleaning, cute, shy
Table 5: Selected stereotypes for discussion and their associated demographic groups in Llama 3 8B.
5.2 Muslim Terrorist Stereotypes
Arab Male given names are disproportionately found to have a low perplexity for the words extremist and terrorist. Research has found a common narrative of all terrorists being Muslim, and sometimes this narrative even being extended to suggest that all Muslims are terrorists Chang and Kleiner (2003); Corbin (2017). This association also has drawn criticism from media scholars, arguing that such portrayals demonize and dehumanize Arab individuals, portraying them as brutal religious extremists Shaheen (2003); Najm (2019). This stereotype has recently been found to be more prevalent in AI generated content than human generated content Narayanan Venkit et al. (2023).
5.3 Intersectional Harms
Recent work states that "researchers overwhelmingly reduce intersectionality to optimizing for fairness metrics over demographic subgroups." Ovalle et al. (2023). Although we look at demographic subgroups within this work, we also note the importance of discussing the power relations and social contexts in which these biases exist, and for which groups they are most likely to cause harm.
One such bias is the continuing and damaging perception of Asian women as docile and submissive Zheng (2016); Azhar et al. (2021). Table 5 shows descriptors cute and shy associated with Japanese women and quiet associated with Chinese women. The stereotype of Japanese women as shy reflects an Orientalist view of Japan, and may also reflect the disadvantaged social position in which Japanese women in the West are situated rather than any essential commonality among them Kitamura (2005). This reflects the context in which many of the LLMs tested have been trained - on Internet data over-representing the West Bender et al. (2021).
Lai (1992) discusses the continuing perception of Asian women as "cute (as in doll-like), quiet rather than militant, and unassuming rather than assertive". The nature of these characterizations speaks to a lack of respect afforded to Asian women as self-sufficient, complex individuals Matsumoto (2020), and contributes to the development of internalized racism and sexism Museus and Truong (2013).
Further, consider the stereotype of Asian Americans as “good at math”. This reinforces subordination along the lines of identity by dictating how Asian Americans and other minorities are expected to behave, and disregards the experiences of Asian Americans who do not achieve model minority success, potentially impacting their self-worth Lee (1999). Such stereotypes perpetuate harmful biases and reinforce societal inequalities.
6 Stereotypes via Generation
The above experiment sheds light on the plausibility assigned to sentences by LLMs containing combinations of proper nouns and descriptors. We complement this experiment by directly looking at models’ generations, which has advantages such as potentially higher correlation with downstream performance Luden et al. (2024). To this end, we use a zero-shot prompting method that utilizes the given names in GlobalBias. Our prompt (Appendix B) instructs the model to generate a dataset of characters, each associated with a given name from GlobalBias, with information such as hobbies, personality traits and physical attributes. An example can be found in Table 1. Additionally, the prompt instructs the model to ensure that the dataset is free from stereotypes and clichés, and to treat all names equally. Our experiment encompassed four models with widespread usage: Claude 3 Opus, Llama 3 70B Instruct, and OpenAI’s GPT 3.5 and GPT 4o. [^2] We use a temperature of 1 to ensure a wide variety of outputs. The outputs were generated 3 times for each model, resulting in 1200 character profiles for each model.
The rationale for using an open-ended generation setting was two-fold: (1) the likelihoods studied in the previous section do not always correspond to model outputs Parrish et al. (2022), and (2) taking a lexicon-free approach allows us to capture stereotypes that we had not thought of a priori. Furthermore, this approach enables testing for stereotypes in closed-source models.
6.1 Classification
To assess the the level of bias in each model, we construct a one-vs-all SVM classification across gender, ethnicity, and gender-by-ethnicity groups, to measure how easily differentiable demographic groups are from each other. We partition our data in to 70% for training and 30% for testing, stratified based on demographic group. Each character profile was represented using 11 features, with each feature encoded as either a one-hot vector (for single words) or sparse vector of the relative frequencies of the words in the feature (for lists of words).
Our results show that character descriptions corresponding to different demographic names are distinguishable from one another by gender, ethnicity and the intersection of the two, indicating that all four models produce stereotypical outputs, even when explicitly instructed not to (Table 6).
Notably, GPT-4o exhibits the highest level of distinction between groups. The SVM achieved an accuracy of 33.3%, over 13 times higher than a baseline accuracy of random classification (2.5%) which would indicate no difference between demographic groups. Previous research has demonstrated that larger models tend to exhibit greater gender and racial biases Ganguli et al. (2022); Rae et al. (2022); Ganguli et al. (2023). Our study extends these findings by revealing that this pattern also manifests in intersectional groups in the context of stereotypes.
Model Gender + Ethnicity Ethncity Gender
Llama 3 70B 18.3% 30.6% 83.3%
GPT 3.5 21.7% 32.2% 88.9%
Claude 3 Opus 26.4% 36.1% 91.9%
GPT 4o 33.3% 38.6% 93.9%
Table 6: SVM classification accuracy for character profiles of different demographic groups. A lower accuracy indicates more similar character profiles across groups, therefore less stereotypical outputs. The task involved classification of 40 groups for Gender + Ethnicity accuracy, 20 groups for Ethnicity and 2 for Gender.
6.2 Feature Analysis
We conduct a feature elimination process to identify the importance of different features in distinguishing between demographic groups, in order to identify potential sources of bias. We analyse groups of features such as ’hobbies’, rather than individual features such as ’reading’. A full table of the impact of each group of features can be found in Appendix C.
We find that, across all models, religion is the most influential feature in predicting ethnicity. For 3 out of 4 models, religion is also the strongest feature when classifying combined gender and ethnicity groups suggesting that models are overly reliant on religious features when describing ethnicity, potentially leading to biased or inaccurate portrayals of individuals. Conversely, for predicting gender alone, removing religion from the input results in increased accuracy. Similarly, skin colour is a significant feature for ethnicity and gender + ethnicity classifications, while it has minimal impact on gender-only. Significant features that emerged for gender-only classification were physical characteristics such as height and build.
Our results also show that combining features from gender-only and ethnicity-only classifications does not lead to improved performance in gender + ethnicity groups. For example, in Claude 3 Opus, the inclusion of sexual orientation decreased accuracy in ethnicity-only and no effect in gender-only classifications, while improving accuracy in gender + ethnicity classification. This highlights that intersectional identities and the stereotypes that affect them are more complex than the sum of their parts Crenshaw (1989), and underscores the significance of considering intersectionality when evaluating bias to foster fair and inclusive AI systems.
6.3 Top Words
Building on the ranking of individual features, we use Jensen-Shannon divergence (JSD) to identify differentiating words for each gender-by-ethnicity group across different features Trujillo et al. (2021); Cheng et al. (2023). We utilize the Shifterator implementation of JSD Gallagher et al. (2021) to compute the top 10 words for each feature, and the groups they belong to. The top words for selected features for Llama 3 70B Instruct and GPT 40 (best and worst models) can be found in Appendix D.
Given that religion emerged as the most significant feature for both gender-by-ethnicity and ethnicity-only groups in our analysis, we examine it further here. As illustrated in Table 7, the top religions identified by JSD and the gender-by-ethnicity groups for which they were generated align consistently with the groups they were correlated with via APX, demonstrating that bias stays consistent across the model’s internal representations and generative outputs, in contrast to claims made in Parrish et al. (2022).
The association of certain religions with demographic groups reinforces essentializing narratives, such as the conflation of the Islamic world and the Arab world Chang and Kleiner (2003). Instead of representing the diversity within groups, the perpetuation of religious stereotypes defines each of these demographic groups solely based on a limited, fixed set of characteristics—such as being Muslim or from the Middle East—rather than recognizing their full humanity Rosenblum and Travis (1996); Woodward (1997). The persistence of religious stereotypes in LLM outputs may further marginalize individuals from other religious and geographic backgrounds with certain given names.
Word Generation APX
jewish Israeli, M Israeli, F Israeli, M Israeli, F
hindu Indian, M Indian, F Indian, M Indian, F
shinto Japanese, M Japanese, F Japanese, M Japanese, F
buddhist Thai, M Thai, F Thai, M Thai, F
muslim Arab, M Turkish, M Arab, M Arab, F
Table 7: Top differentiating religion words and associated groups in both experiments using Llama 3 70B (Generation) and Llama 3 8B (APX).
7 Conclusion
In this work, we present the GlobalBias dataset, which allows us to undertake a comprehensive study of intersectional stereotypes. We introduce a new evaluation metric, APX, to adjust for high-frequency given names in training. This study examines a broader range of demographic groups than previous studies, and we conduct multiple experiments that investigate both the model’s internal representations via APX and model outputs via generation experiments.
We find that larger models produce more stereotypical outputs, even when explicitly instructed not to. We also show using the example of religion that bias stays consistent across model’s internal representation and outputs. Through this investigation, we aim to raise awareness about the importance of considering intersectionality when evaluating models and encourage nuanced and thoughtful evaluation of stereotypes in LLMs.
Limitations
While our work aims to broaden the scope of ethnicities covered in NLP bias research, there are many ethnic groups and genders not covered in this work, and we exclude other critical aspects such as age, disability, and socioeconomic status. The dataset’s creation process excludes gender-neutral names, limiting its applicability to a broader spectrum of identities. Moreover, the GlobalBias dataset is not intended as a benchmark; instead, it is used to gain insights into a wider set of intersectional demographic groups.
By explicitly categorizing and associating stereotypes with specific demographic groups, there is a risk of perpetuating the very biases the study aims to mitigate. The study does not propose specific debiasing techniques, and while the GlobalBias dataset and APX metric can aid future efforts, practical implementations and evaluations of debiasing strategies are needed.
Furthermore, other measures for perplexity have been proposed such as AULA Kaneko and Bollegala (2021). We use perplexity, and APX, as it can be adapted for use across a range of model architectures. The evaluation methods, while insightful, may not fully reflect real-world scenarios. Findings, particularly regarding larger models producing more stereotypical outputs, are based on current LLM architectures and may need re-evaluation as new models emerge. The closed-source nature of some models also limits transparency and replicability.
Acknowledgements
We would like to thank Nedjma Ousidhoum and Yi Zhou for their very helpful comments in reviewing this paper. We also thank Dimosthenis Antypas, Joanne Boisson, Jose Camacho-Collados and Hsuvas Borkakoty for helpful feedback. This work is funded in part by the UKRI AIMLAC CDT.
Appendix A Stereotypes via APX
Group Associated Descriptors
African Female privileged, radical, ugly, underserved, very loud
Arab Female Muslim, refugee
Arab Male disabled, extremist, moderate, Muslim, terrorist
Baltic Male bald-headed
Chinese Female Confucianist, good at distance running, good at math, quiet, small, Taoist, very short, very smart, white-collar
Chinese Male dishonest, larger, small, smaller, Taoist, white-collar
Dutch Female hetero, mildly overweight, slim, sweaty
Dutch Male hetero, hipster
English Female AFAB, All-American, AMAB, controlling, elite, Gen-X, little person, morbidly obese, Mormon, nondisabled, pro-choice, pro-life
English Male addict, All-American, bully, can’t get laid, child predator, conspiracy theorist, country, Evangelical, felon, huge, massive, Mormon, murderer, pro-life, stutterer
French Male sedentary, snob
Greek Male athletic, Christian, scientist
Hispanic Female brown-skinned, Jehovah’s Witness, pale-skinned, transfeminine, transsexual
Hispanic Male macho
Hungarian Male hungry, hunky
Indian Female affluent, ambitious, differently abled, engaged, expressive, fit, Hindu, homemaker, illiterate, independent, Jain, married, paralyzed, patient, progressive, shorter, Sikh, single, spiritual, supportive, underprivileged, underweight, uneducated, unmarried, urban, well educated
Indian Male atheist, differently abled, dishonest, fit, gangly, Hindu, illiterate, Jain, moneyed, paralyzed, patient, Sikh, spiritual, Spiritualist, supportive, taller, uneducated, unemployed, unmarried, well-built
Israeli Female Druze, Hasidic, irreligious, Jew, Jewish
Israeli Male complaining, Conservative, convert, Druze, goy, Hasid, Hasidic, irreligious, Jew, Jewish, religious, right-wing
Japanese Female always cleaning, cute, genderless, happy, have no friends, JAP, ponytailed, scary, Shintoist, shy, third-gender, very short
Japanese Male macho, scary, Shintoist
Korean Female acne-covered, Cheondoist, skinny
Korean Male Cheondoist, lack social skills, underemployed, unhygenic
Nordic Male demiboy, tall
Slav Male physically fit
Thai Female Buddhist
Thai Male bad driver, Buddhist, effeminate, handicapped, villager
Turkish Female refugee
Turkish Male asylum seeker, refugee, wear a turban
Table 8: All stereotypes and their associated demographic groups from Llama 3 8B, measured taking a 1% one-tailed significance level of our bias score across the 40 demographic groups for each descriptor.
Appendix B Generation Prompt
I am creating a dataset of different characters for an educational game. The dataset must be in JSON with the fields:
"name" - string
"age" - integer
"personality_traits" - list of strings, length 3
"negative_traits" - list of strings, length 3
“hobbies” - list of strings, length 3
"occupation" - string
"physical_characteristics" - dictionary with keys height (in ft), hair_colour, eye_colour, skin_colour, build
"special_move" - string
“socioeconomic_status” - string
“sexual_orientation” - string
“religion” - string
Could you create me profiles for characters named: { list_of_names }
I would like my dataset to be unbiased and avoid clichés. Please treat all names the same and avoid using stereotypes.
Please output only JSON with no additional text.
Appendix C Feature Analysis in Generation
Feature Eliminated Llama 3 70B GPT 3.5 Claude 3 Opus GPT 4o
Overall Accuracy (%) 18.3% 21.7% 26.4% 33.3%
religion -4.1% $\downarrow$ -4.8% $\downarrow$ -3.6% $\downarrow$ -8.0% $\downarrow$
hair_colour -1.1% $\downarrow$ -0.3% $\downarrow$ -0.3% $\downarrow$ -1.1% $\downarrow$
height -0.2% $\downarrow$ -2.0% $\downarrow$ -3.6% $\downarrow$ -3.6% $\downarrow$
sexual_orientation +0.3% $\uparrow$ 0.0% -0.6% $\downarrow$ +1.7% $\uparrow$
hobbies +0.9% $\uparrow$ -0.6% $\downarrow$ +0.5% $\uparrow$ -1.4% $\downarrow$
build +0.9% $\uparrow$ -0.6% $\downarrow$ -0.6% $\downarrow$ -1.4% $\downarrow$
socioeconomic_status +2.3% $\uparrow$ +1.1% $\uparrow$ +1.4% $\uparrow$ +0.6% $\uparrow$
skin_colour +0.6% $\uparrow$ -1.1% $\downarrow$ -3.1% $\downarrow$ -2.2% $\downarrow$
eye_colour +2.8% $\uparrow$ +0.5% $\uparrow$ +2.2% $\uparrow$ +1.7% $\uparrow$
personality_traits +1.4% $\uparrow$ -1.7% $\downarrow$ +0.8% $\uparrow$ -0.2% $\downarrow$
negative_traits +2.0% $\uparrow$ -1.7% $\downarrow$ +2.2% $\uparrow$ 0.0%
age +2.8% $\uparrow$ +0.5% $\uparrow$ +0.8% $\uparrow$ +0.6% $\uparrow$
occupation +0.6% $\uparrow$ -0.3% $\downarrow$ +0.3% $\uparrow$ -1.4% $\downarrow$
Table 9: Model accuracies and feature impact on differentiation accuracy across demographic groups. The arrows indicate whether the feature caused the accuracy to go up (green) or down (red), with the change in accuracy shown.
Feature Eliminated Llama 3 70B GPT 3.5 Claude 3 Opus GPT 4o
Overall Accuracy (%) 30.6 32.2 36.1 38.6
religion -11.4% $\downarrow$ -6.4% $\downarrow$ -7.8% $\downarrow$ -11.4% $\downarrow$
eye_colour -2.5% $\downarrow$ -4.4% $\downarrow$ -1.1% $\downarrow$ -1.9% $\downarrow$
skin_colour -0.3% $\downarrow$ -5.0% $\downarrow$ -7.5% $\downarrow$ -4.7% $\downarrow$
negative_traits 0.0% -2.5% $\downarrow$ +0.6% $\uparrow$ +0.3% $\uparrow$
personality_traits +0.2% $\uparrow$ -2.5% $\downarrow$ +2.0% $\uparrow$ -0.3% $\downarrow$
build +0.2% $\uparrow$ -3.0% $\downarrow$ +1.7% $\uparrow$ -0.3% $\downarrow$
occupation +0.8% $\uparrow$ -2.2% $\downarrow$ +0.6% $\uparrow$ -0.5% $\downarrow$
hobbies +1.1% $\uparrow$ -1.4% $\downarrow$ +0.3% $\uparrow$ -0.5% $\downarrow$
sexual_orientation +1.1% $\uparrow$ 0.0% +2.2% $\uparrow$ -1.1% $\downarrow$
socioeconomic_status +1.3% $\uparrow$ -0.5% $\downarrow$ +2.0% $\uparrow$ -0.5% $\downarrow$
height +1.6% $\uparrow$ -0.5% $\downarrow$ -1.1% $\downarrow$ -1.4% $\downarrow$
hair_colour +2.2% $\uparrow$ -3.9% $\downarrow$ -1.1% $\downarrow$ -0.5% $\downarrow$
age +2.5% $\uparrow$ -1.9% $\downarrow$ +1.4% $\uparrow$ -1.1% $\downarrow$
Table 10: Model accuracies and feature impact on differentiation accuracy across ethnicities. The arrows indicate whether the feature caused the accuracy to go up (green) or down (red), with the change in accuracy shown.
Feature Eliminated Llama 3 70B GPT 3.5 Claude 3 Opus GPT 4o
Overall Accuracy (%) 83.3 88.9 91.9 93.9
height -4.7% $\downarrow$ -7.0% $\downarrow$ -7.7% $\downarrow$ -10.0% $\downarrow$
negative_traits -1.6% $\downarrow$ +0.5% $\uparrow$ -0.2% $\downarrow$ -0.6% $\downarrow$
hair_colour -1.4% $\downarrow$ -0.3% $\downarrow$ -1.3% $\downarrow$ -0.6% $\downarrow$
eye_colour -1.1% $\downarrow$ -0.6% $\downarrow$ 0.0% -0.3% $\downarrow$
occupation -0.8% $\downarrow$ +0.5% $\uparrow$ 0.0% 0.0%
age -0.5% $\downarrow$ 0.0% +0.9% $\uparrow$ -0.8% $\downarrow$
hobbies -0.5% $\downarrow$ -0.6% $\downarrow$ -0.5% $\downarrow$ -1.1% $\downarrow$
religion -0.5% $\downarrow$ +1.4% $\uparrow$ -0.5% $\downarrow$ +0.5% $\uparrow$
personality_traits 0.0% 0.0% -0.2% $\downarrow$ -0.6% $\downarrow$
sexual_orientation 0.0% -0.3% $\downarrow$ 0.0% -2.5% $\downarrow$
socioeconomic_status +0.3% $\uparrow$ +1.1% $\uparrow$ +0.3% $\uparrow$ -0.3% $\downarrow$
skin_colour +0.3% $\uparrow$ +0.3% $\uparrow$ -0.2% $\downarrow$ -0.3% $\downarrow$
build +1.7% $\uparrow$ -0.8% $\downarrow$ -1.9% $\downarrow$ -1.1% $\downarrow$
Table 11: Model accuracies and feature impact on differentiation accuracy across gender. The arrows indicate whether the feature caused the accuracy to go up (green) or down (red), with the change in accuracy shown.
Appendix D Top Words in Generation
Feature Word Associated Groups
negative_traits arrogant manipulative perfectionist pessimistic selfish Baltic Female, English Male Japanese Female, Chinese Female, Baltic Female Slav Female, French Male English Male, African Male Israeli Male
hobbies yoga painting dancing playing piano Indian Female, Thai Female, African Male, Arab Male, English Male, Baltic Male Arab Female, African Male Nordic Female Chinese Female
occupation politician rabbi freelance writer social worker therapist event planner nurse engineer counselor software engineer Turkish Male Israeli Male German Female Arab Female French Female Israeli Female African Female Arab Male Israeli Female Korean Male
socioeconomic_status upper middle class lower class upper class lower middle class working class middle class African Female Nordic Male, Hispanic Male Baltic Female, Greek Female, Indian Female, Greek Male English Female Italian Female Israeli Male
sexual_orientation bisexual pansexual asexual homosexual Greek Male, Hispanic Male, Hispanic Female, German Male, Hungarian Male French Female, Indian Male, Israeli Male Japanese Male English Female
religion jewish hindu shinto buddhist muslim Israeli Male, Israeli Female Indian Male, Indian Female Japanese Female, Japanese Male Thai Female, Thai Male Arab Male, Turkish Male
hair_colour black dark brown curly brown English Female, Baltic Male, Italian Female, Dutch Male, Slav Male, African Female, Nordic Male German Female Greek Female
skin_colour fair dark Arab Male, African Male, Thai Male, Indian Male French Female, Indian Female, Baltic Female, African Female, Italian Female, Baltic Male
Table 12: Top 10 differentiating words across all groups for selected features in Llama 3 70B Instruct.
Feature Word Associated Groups
negative_traits shy impulsive aloof stubborn disorganized stern rigid perfectionist overcritical Japanese Female Japanese Male, Italian Male Slav Male English Male Nordic Female German Male German Male Slav Female Thai Female
hobbies calligraphy cycling painting yoga cooking soccer origami Chinese Female, Japanese Male, Chinese Male Dutch Male Italian Female French Female Thai Female, Italian Male African Male Japanese Female
occupation chef research scientist data scientist software developer graphic designer historian mechanical engineer professor journalist Thai Female, Italian Male Chinese Male Chinese Male Baltic Male Nordic Female German Male Nordic Male Indian Male Baltic Female
socioeconomic_status middle-income upper middle class middle middle-class Slav Female, German Male Japanese Female, Korean Female, German Female Italian Male, Nordic Male, Greek Male Turkish Male, Indian Male
sexual_orientation lesbian gay asexual bisexual Slav Female, Dutch Male, Turkish Male, Israeli Female French Male, African Female Japanese Female Japanese Male, Italian Female
religion jewish hindu muslim shinto catholic buddhist christian Israeli Female, Israeli Male Indian Female, Indian Male Arab Male Japanese Male Italian Male Thai Female, Thai Male African Female
hair_colour blonde black brown Nordic Female, Turkish Male, Arab Male, Greek Male German Male, Dutch Male, Nordic Male, Baltic Male Italian Female, Japanese Male
skin_colour fair light tan dark olive Thai Female, Arab Male, Thai Male Japanese Male African Female, African Male French Female, English Male, Hungarian Female, Italian Male
Table 13: Top 10 differentiating words across all groups for selected features in GPT 4o.